### v2

# A novel generalized inner product-based wave scattering from an underwater source in a compressible ocean

Abstract
Motivated by applications to underwater explosions and volcanic eruptions, this paper considers the evolution of an initial pressure disturbance in the ocean, including effects due to the dynamic and static compression of water and the free surface. In order to solve the equations of motion of a linear compressible ocean, a special inner product is introduced, which allows us to apply self-adjoint operator theory. What results is a Hilbert space in which the acoustic-gravity modes are orthogonal in the generalised sense. This allows the time-domain evolution of the free surface and subsurface pressure field resulting from an initial disturbance to be calculated. Our simulations show initial radial propagation of the pressure pulse and subsequent reflection from the water surface and the rigid ocean floor, eventually leading to horizontal propagation away from the source point. The solutions with and without the inclusion of the static compression are compared, and the effect of static compression is shown to be small but not negligible.
1 Introduction
The evolution of the pressure field from an initial disturbance in the ocean has been of significant interest to the scientific community for a long time, primarily due to the potential impact of underwater nuclear bomb testing, underwater volcanic eruptions, seismic events, and the associated tsunami hazards. There also exists significant monitoring of pressure; for example, the Comprehensive Nuclear-Test-Ban-Treaty Organization (CTBTO) monitors pressure signals from hydrophone stations to detect possible nuclear tests. Similar devices also exist for detecting tsunami waves. Accurate modelling of the pressure signatures allows information about the event to be determined, for example, an attempt has been made to estimate the crash site location of Malaysian Airlines Flight 370 (MH370) using an inverse problem (Kadri, 2024). The accuracy of such inverse problems depends on the accurate mathematical analysis of the associated wave evolution problem. Obviously, the short–time behaviour of such an event may be highly nonlinear and involves the formation of a cavity and subsequent expansion at the initial state. However, after some time (and some distance), the pressure wave inside the water column can be accurately modelled with a linearised water wave theory that accounts for the effect of dynamic compression (we use this term to distinguish the effect of the local compressibility from the static compression which is the accumulation of compression due to the weight of the overlying water column. In this work, we consider the linear problem and solve the initial value problem, including the effect of the free surface. The solution relies on the use of a special Hilbert space. We also extend this formulation to the more challenging case where we include the effect of static compression.
Many energy-conserving linear wave problems can be recast as an abstract wave equation $\partial_{t}^{2}u+\mathcal{A}u=0$, where $\mathcal{A}$ is a positive self-adjoint operator in a Hilbert space $(\mathcal{H},\langle\cdot,\cdot\rangle)$. For example, choosing $\mathcal{A}=-c^{2}\partial_{x}^{2}$ gives the classical wave equation with wave speed $c$. More complicated forms of $\mathcal{A}$ appear in various scenarios such as hydroelasticity (Meylan, 2002). While the use of the spectral theorem to obtain the solution as an expansion over the eigenfunctions of $\mathcal{A}$ was long known when $\mathcal{A}$ has a discrete spectrum (Fourier/Sturm Liouville) the case for when $\mathcal{A}$ has a continuous spectrum began being worked out for the Schrödinger equation with the advent of quantum mechanics (Povzner, 1953; Ikebe, 1960). The theory, which has been extended to the classical wave equation (Wilcox, 1975) and to hydrodynamics for the case of incompressible water wave scattering (Hazard & Loret, 2007; Hazard & Meylan, 2008), is known as a generalised eigenfunction expansion. The theoretical foundation is given by a Gelfand triple, also known as a rigged Hilbert space, which is a mathematical structure used to extend the usual Hilbert space framework to include more generalised functions (like distributions). This is because we need to deal with eigenfunctions that are not square-integrable (i.e., not in the Hilbert space) but are still useful and meaningful in a generalised sense. In general, this theory has been developed in isolation from computation, despite the fact that the simplest example of this structure is the Fourier transform, and its computation by the discrete Fourier transform (often implemented using the fast Fourier transform) is one of the most ubiquitous algorithms.
It turns out that the theory of generalised eigenfunctions and spectral solutions for continuous spectra can be obtained numerically by applying quadrature to the integrals over the continuous spectrum. This idea was in the work of Hazard & Meylan (2008); Meylan (2009); Meylan & Eatock Taylor (2009); Peter & Meylan (2010), but appears explicitly in Wilks et al. (2024, 2025). As far as we are aware, this work is the first application of the theory to compressible water wave scattering. A key advantage of the numerical spectral technique is that the generalised eigenfunctions, which satisfy $\mathcal{A}\hat{u}=-\omega^{2}\hat{u}$, are precisely the time harmonic solutions for the angular frequency $\omega$. These single-frequency scattering solutions can be found numerically for a wide range of problems, e.g. variable bathymetry, which would be a relatively straightforward extension of the present work. We note that the ease of determining the spectrum stands in contrast with a more numerical method, in which the operator $\mathcal{A}$ is approximated by a matrix and the spectrum of this matrix is sought. In this case, very many numerical problems arise, especially because the eigenvalues of the underlying operator are continuous. Further problems can arise from the fact that it may be difficult to preserve the self-adjoint property of the matrix if the inner product is nontrivial (as in the case with the current problem).
The inclusion of the compressibility of water is an essential part of our work. The introduction of realistic but slight water compressibility produces acoustic-gravity waves (AGW) - hydroacoustic waves influenced by gravity at low frequency, which act as analogous propagating modes in acoustic/electromagnetic waveguides. These modes capture the effect of the boundary condition of the free surface and play a pivotal role in the accurate calculation of the pressure of the water column. It is worth mentioning that at the limit of infinite sound speed in the water, the first AGW mode transforms into the incompressible wave mode under realistic parameter choices. Following the devastating aftermath of the $2004$ Indian Ocean tsunami, a renewed interest in the application of AGW in tsunami detection mechanisms (through modelling of surface wave propagation) has grown through the works of (Nosov, 1999; Nosov & Skachko, 2001; Nosov & Kolesov, 2007), which are based on the well-established fundamental mathematical theory developed by Longuet-Higgins (1950), Sells (1965), Miyoshi (1954), Yamamoto (1982). While the incompressible water wave theory is the most common in practice, the compressibility theory pointed out its efficiency through an enhanced accuracy in estimating the observed tsunami wave profile. Following the development, Stiassnie (2010) and Hendin & Stiassnie (2013) computed the far-field profiles of the tsunami and AGW modes. Subsequently, there has been a surge in understanding different properties of these AGW modes (e.g., two interacting gravity waves producing AGW(Kadri & Stiassnie, 2013), triad resonances (Kadri, 2016; Kadri & Akylas, 2016)), impact on different oceanic processes (deep-water transport (Kadri, 2014), ice shelf breaking (Kadri & Stiassnie, 2012), wave dissipation on sea ice (Chen et al., 2019)), impact of seafloor characteristics (porous (Chierici et al., 2010) and viscous (Abdolali et al., 2015)). Different mathematical techniques are also deployed to solve the associated initial boundary value problem (IBVP) (rigid but variable bathymetry using depth-integrated mild-slope equations (Sammarco et al., 2013), Fourier-series expansion (Das & Meylan, 2023b), eigenfunction expansion (Das & Meylan, 2023a)), which are essential to address the static compression, an outcome of the water density variation along the depth due to its own weight.
The static compression of the ocean is due to the effect of the slight compressibility of water, causing an increase in the density of the deep ocean. It is only significant for waves which penetrate the deep ocean, and even then, its effect is small. The initial account of static compression can be found through the pioneering work of Longuet-Higgins (1950). However, this quantity is rarely included in the mathematical formulation, citing its presumably insignificant impact on the wave propagation until the work of Abdolali & Kirby (2017), who illustrated the amplification of tsunami phase speed at a higher ocean depth induced by static compression. A few contemporary works retained static compression in the mathematical treatment of physical problems to study the behaviour of the pure-gravity mode (Kadri, 2015; Abdolali & Kirby, 2017; Abdolali et al., 2019; Kadri, 2016; Kadri & Stiassnie, 2013). The most notable one is by Omira et al. (2022) to match the Hunga Tonga–Hunga Ha‘apai (HTHH) volcanic eruption-induced tsunami data using the resonance between the acoustic waves and the surface gravity wave. The quantification of the effect of static compression is entangled with the varying sound speed along the water depth, and the above literature makes use of the Eulerian approach of solution based on the approximation of the vertical sound profile. Michele & Renzi (2020) used a perturbation technique to derive an analytical expression correct up to third order that governs the evolution of the hydroacoustic waves. Recently, Andrade et al. (2025) used a dynamic approach to study triad resonances and energy exchange between different modes. Recent developments include a novel Lagrangian method to approach the density stratification through linearization of the compressible Euler equation (Dubois et al., 2023), and it has been demonstrated that other models in the literature can be derived from their general proposed model. In the context of tsunami wave generation, Das & Meylan (2023a) solved for the time-dependent evolution due to impulsive forcing from the sea floor, including the effect of static compression. The eigenfunction matching required a special inner product, which is generalised here. This novel method was later extended to include more realistic geometries (a step (Das et al., 2024), a trench (Pethiyagoda et al., 2024b), a generic ocean floor (Pethiyagoda et al., 2025a)), three-dimensional solutions (Das & Meylan, 2024), and surface disturbance-induced wave propagation models (Pethiyagoda et al., 2025b, 2024a) that augment the works where static compression is dropped (Renzi & Dias, 2014) and ocean compressibility is ignored (Liu & Higuera, 2022). We note here that there are ways in which the current work could be extended to include other effects which are non-standard beyond the static compression we consider here. Two examples are the effect of the bottom elasticity (Eyov et al., 2013; Abdolali et al., 2019; Kadri, 2019; Williams & Kadri, 2023) and more complex static compression using sound speed vertical profile (Michele & Renzi, 2020). We note that, for the range of parameters we investigate here the effect of the static compression is small, even for the simplest formulation which we use.
The properties of the operator involved in this class of problems, where ocean water compressibility (including static compression) is involved, are not well studied in the literature. We provide a mathematical structure for the operators involved in a rigid and flat ocean floor via a special inner product. The manuscript is structured as follows. Section 2 presents the two-dimensional mathematical formulation of the physical problem where static compression is ignored, followed by the solution method that involves deriving a generic two-dimensional inner product in §3. The time domain simulations are detailed in §4. The problem is much simpler if static compression is ignored, and for many scenarios, only this solution will be required, and this is the motivation for giving this in detail. In §5, we extend the formulation to static compression and calculate the time-domain simulations of the flow field in this case. We note that the inclusion of static compression complicates the derivation, but the numerical solution takes a similar computational time to the solution without compression. The manuscript concludes with a brief summary in §6.
2 Mathematical formulation without static compression
We construct the physical problem in a homogeneous ocean water environment where the standard density variance due to the water column’s own weight is ignored, leading to ocean compressibility without static compression (only the dynamic compression is included). The mathematical problem is formulated in a two-dimensional coordinate system $(x,z)$ whose $x$-direction is horizontal and $z$-direction vertically upwards. Linearised water wave theory is used to model the wave scattering induced by an initial pressure in a compressible ocean. The water is considered inviscid and the motion is irrotational, leading to the existence of a time-dependent velocity potential $\Phi(x,z,t)$ throughout the fluid domain such that $\nabla\Phi(x,z,t)=(u(x,z,t),\,v(x,z,t))$ where $u(x,z,t)$ and $v(x,z,t)$ are the fluid velocities in the $x$ and $z$ directions, respectively. The equilibrium free surface and the rigid ocean floor are defined by $z=0$ and z=hz=-h, respectively.
The IBVP, as given by (Yamamoto, 1982; Stiassnie, 2010), is
$\Phi_{tt}$ $=c^{2}\nabla^{2}\Phi,\quad$ $(x,z)\in\Omega,$ (1a)
$\Phi_{tt}+g\Phi_{z}$ $=0,\quad$ $z=0,$ (1b)
$\Phi_{n}$ $=0,$ $z=-h,$ (1c)
$\Phi_{x}$ $\to 0,$ $|x|\to\infty,$ (1d)
subject to the initial conditions
$\Phi(x,z,0)$ $=\Phi_{0}(x,z),$ (1e)
$\Phi_{t}(x,z,0)$ $=-\frac{P_{0}(x,z)}{\rho},$ (1f)
where $c$ represents the speed of sound in sea water, $g$ is acceleration due to gravity, $\Omega=\{(x,z):|x|<\infty,-h<z<0\}$ is the fluid domain, and $\Phi_{0}$ and $P_{0}$ are the initial potential and hydrodynamic pressure of the fluid, respectively.
z=0z=0z=hz=-hRigid impermeable bottomxxzz
Figure 1: Schematic diagram of the physical problem showing wave propagation in a compressible ocean (without including static compression) due to an underwater explosion (red star). Pressure waves generated by the explosion (red dashed lines) are reflected off the sea bed and free surface (blue dashed lines) and induce propagating waves on the fluid surface (solid red lines)
We consider this problem in the context of self-adjoint operator theory. We write
$\Phi_{tt}=c^{2}\nabla^{2}\Phi,$ (2)
and find an inner product with respect to which the operator $c^{2}\nabla^{2}\Phi$ is self-adjoint subject to the boundary conditions. To find this inner product, we use Green’s second identity in the following form:
$\iint_{\Omega}\left(\nabla^{2}\Phi\,\Psi-\Phi\,\nabla^{2}\Psi\right){\rm d}V$ $=\oint_{\partial\Omega}\left(\Phi_{n}\,\Psi-\Phi\,\Psi_{n}\right){\rm d}S,$ (3a)
where $\Phi$ and $\Psi$ are the two solutions of the BVP (1a)-(1d). After rearranging the terms, we get
$\iint_{\Omega}\nabla^{2}\Phi\,\Psi{\rm d}V-\oint_{\partial\Omega}\Phi_{n}\,\Psi{\rm d}S=\iint_{\Omega}\Phi\,\nabla^{2}\Psi{\rm d}V-\oint_{\partial\Omega}\Phi\,\Psi_{n}{\rm d}S.$ (3b)
Taking $\Omega$ to be the region $(-L,L)\times(-h,0)$ gives, in the limit as $L\to\infty$,
$\iint_{\Omega}\nabla^{2}\Phi\,\Psi{\rm d}V-\int_{-\infty}^{\infty}\Phi_{z}\,\Psi\,{\rm d}x-\lim_{x\rightarrow\infty}\int_{-h}^{0}\Phi_{x}\,\Psi\,{\rm d}z+\lim_{x\rightarrow-\infty}\int_{-h}^{0}\Phi_{x}\,\Psi\,{\rm d}z$
$=\iint_{\Omega}\nabla^{2}\Psi\,\Phi{\rm d}V-\int_{-\infty}^{\infty}\Psi_{z}\,\Phi\,{\rm d}x-\lim_{x\rightarrow\infty}\int_{-h}^{0}\Psi_{x}\,\Phi\,{\rm d}z+\lim_{x\rightarrow-\infty}\int_{-h}^{0}\Psi_{x}\,\Phi\,{\rm d}z,$ (3c)
where we have used (1c). Finally, use of (1d) gives
$\iint_{\Omega}\nabla^{2}\Phi\,\Psi{\rm d}V+\frac{c^{2}}{g}\int_{-\infty}^{\infty}\nabla^{2}\Phi\,\Psi\,{\rm d}x=\iint_{\Omega}\Phi\,\nabla^{2}\Psi{\rm d}V+\frac{c^{2}}{g}\int_{-\infty}^{\infty}\Phi\,\nabla^{2}\Psi\,{\rm d}x.$ (3d)
Here we used the condition on $\Phi$ and $\Psi$ and a combination of (1a) and (1b) that leads to $\Phi_{z}=-\frac{c^{2}}{g}\nabla^{2}\Phi$.
We define the Hilbert space to be the space equipped with the inner product
$\langle\Phi,\Psi\rangle_{\mathcal{E}}=\int_{\Omega}\Phi(x,y)\overline{\Psi(x,y)}\,{\rm d}x{\rm d}y+\frac{c^{2}}{g}\int_{-\infty}^{\infty}\Phi(x,0)\overline{\Psi(x,0)}\,{\rm d}x.$ (4)
for $\Phi$, $\Psi\in\mathcal{E}$. Self-adjoint operator theory requires an inner product space in which the operator has a symmetry, as we have just shown, and also that the domain of the operator and its adjoint are the same. The latter point is rather subtle, but in our case, this will follow when the space is closed. It is worth noting that the operator and the Hilbert space are closely related and that when we close the functions, this closure occurs in the norm defined by our inner product. In what follows, we use the inner product to calculate all projections, which is critical to the solution we develop.
3 Solution technique
Since the operator $c^{2}\nabla^{2}$ is assumed to be self-adjoint, we know that its spectrum is real. In fact, it has a continuous spectrum occupying the entire positive real line. To find the generalised eigenfunctions, we solve
$c^{2}\nabla^{2}\phi$ $=-\omega^{2}\phi,$ $(x,z)\in\Omega,$ (5a)
$-\omega^{2}\phi+g\phi_{z}$ $=0,$ $z=0,$ (5b)
$\phi_{n}$ $=0,$ $z=-h,$ (5c)
subject to an incident wave condition. Solutions of (5a) are nothing more than the frequency domain solutions found by assuming time harmonicity, namely
$\Phi(x,z,t)=\mathrm{Re}\{\phi(x,z)\mathrm{e}^{-{\rm i}\omega t}\},$ (6)
where $\omega$ is the angular frequency.
Using separation of variables, we obtain explicit expressions for the solutions of (5a) of the form
$\phi^{\pm}_{n}(x,z;k)=\mathrm{e}^{{\pm\rm i}kx}f_{n}(z;k),$ (7)
for $n\in\mathbb{N}$ and $k\in\mathbb{R}$, where the depth dependent functions satisfy the eigenvalue problem
$f_{n}^{\prime\prime}(z;k)=\mu_{n}^{2}f_{n}(z;k),$ (8a)
subject to the boundary conditions
$gf_{n}^{\prime}(0;k)$ $=c^{2}(k^{2}-\mu_{n}^{2})f_{n}(0;k),$ (8b)
$f_{n}^{\prime}(-h;k)$ $=0,$ (8c)
where the prime denotes differentiation with respect to $z$. Equation (8) has a discrete spectrum of eigenvalues $\mu_{n}$, indexed by $n\in\mathbb{N}_{0}$, satisfying the following dispersion relation
$\mu_{n}(k)\tanh{\mu_{n}(k)h}+\frac{c^{2}}{g^{2}}\mu_{n}(k)^{2}=\frac{c^{2}k^{2}}{g^{2}},$ (9)
with $\mu_{0}(k)$ being the positive real root and $\mu_{n}(k)$ being of the form $\rm i\tilde{\mu}_{n}(k)$ for $n\geq 1$, where $\tilde{\mu}_{n}(k)$ are real, positive and in ascending order with respect to the index $n$. Note that while $\mu_{n}(k)$ depends on both $n$ and $k$, we will omit the argument $k$ when this is clear from context. The associated vertical eigenfunctions are written as
$f_{n}(z;k)=$ $\frac{\displaystyle\cosh{\mu_{n}(z+h)}}{\cosh{(\mu_{n}h)}}.$ (10)
With reference to (5a), the time harmonic angular frequency associated with the mode $\phi_{n}(\cdot,\cdot;k)$ is
$\omega_{n}(k)=c^{2}(k^{2}-\mu_{n}^{2}).$ (11)
Using the spectral theorem, the general solution of the time domain problem (1) can be written as the following series of integrals over the continuous wavenumber
$\Phi(x,z,t)=\sum_{n=0}^{\infty}\int_{0}^{\infty}\left(C_{n}^{-}(t,k)\phi_{n}^{-}(x,z;k){\rm d}k+C_{n}^{+}(t,k)\phi_{n}^{+}(x,z;k)\right){\rm d}k.$ (12)
Substitution of the above into (1a) then gives
$\sum_{n=0}^{\infty}\int_{0}^{\infty}-\omega_{n}(k)^{2}\left(C_{n}^{-}(t,k)\phi_{n}^{-}(x,z;k){\rm d}k+C_{n}^{+}(t,k)\phi_{n}^{+}(x,z;k)\right){\rm d}k\\ =\sum_{n=0}^{\infty}\int_{0}^{\infty}\left(C_{n}^{-\prime\prime}(t,k)\phi_{n}^{-}(x,z;k){\rm d}k+C_{n}^{+\prime\prime}(t,k)\phi_{n}^{+}(x,z;k)\right){\rm d}k,$ (13)
where the prime denotes differentiation with respect to $t$. Using the following orthogonality results (which we prove in the following section):
[…]
in which $D_{p}(k)$ are normalisation coefficients, $\delta(\cdot)$ is the Dirac delta and $\delta_{np}$ is the Kronecker delta, we obtain
$C_{n}^{\pm\prime\prime}(t,k)=-\omega_{n}(k)^{2}C_{n}^{\pm}(t,k).$ (15)
Solutions to the above ordinary differential equation can be written as
$C_{n}^{\pm}(t,k)=A_{n}^{\pm}(k)\cos(\omega_{n}(k)t)+B_{n}^{\pm}(k)\frac{\sin(\omega_{n}(k)t)}{\omega_{n}(k)},$ (16)
whence
$\Phi(x,z,t)=\sum_{n=0}^{\infty}\int_{0}^{\infty}\left(A_{n}^{-}(k)\phi_{n}^{-}(x,z;k)+A_{n}^{+}(k)\phi_{n}^{+}(x,z;k)\right)\cos(\omega_{n}(k)t){\rm d}k\\ +\sum_{n=0}^{\infty}\int_{0}^{\infty}\left(B_{n}^{-}(k)\phi_{n}^{-}(x,z;k)+B_{n}^{+}(k)\phi_{n}^{+}(x,z;k)\right)\frac{\sin(\omega_{n}(k)t)}{\omega_{n}(k)}{\rm d}k.$ (17)
To obtain the particular solution to the IBVP (1), the coefficient functions $A_{n}^{\pm}$ and $B_{n}^{\pm}$ must be determined from the initial potential and pressure of the fluid, $\Phi_{0}$ and $P_{0}$, respectively. Applying (14) to (17) for $t=0$ yields
[…]
Similarly
[…]
Now we apply the initial conditions (1e) and (1f) and take inner products with respect to $\phi_{p}^{\pm}$ to get
$\langle\Phi_{0}(x,z),\phi_{p}^{\pm}(\cdot,\cdot,k)\rangle_{\mathcal{E}}$ $=\langle\Phi(\cdot,\cdot,0),\phi_{p}^{\pm}(\cdot,\cdot,k)\rangle_{\mathcal{E}}=D_{p}A_{p}^{\pm}(k),$ (19a)
$-\frac{\langle P_{0}(x,z),\phi_{p}^{\pm}(\cdot,\cdot,k)\rangle_{\mathcal{E}}}{\rho}$ $=\langle\Phi_{t}(\cdot,\cdot,0),\phi_{p}^{\pm}(\cdot,\cdot,k)\rangle_{\mathcal{E}}=D_{p}B_{p}^{\pm}(k).$ (19b)
Then the coefficient functions $A_{p}^{\pm}(k)$ and $B_{p}^{\pm}(k)$ are given by
$A_{p}^{\pm}(k)=\frac{\langle\Phi_{0}(x,z),\phi_{p}^{\pm}(\cdot,\cdot,k)\rangle_{\mathcal{E}}}{D_{p}}\quad\mbox{and}\quad B_{p}^{\pm}(k)=-\frac{\langle P_{0}(x,z),\phi_{p}^{\pm}(\cdot,\cdot,k)\rangle_{\mathcal{E}}}{\displaystyle\rho D_{p}}.$ (20)
Hence, we obtain the potential function as
$\Phi(x,z,t)=\sum_{n=0}^{N}\int_{0}^{\infty}\left(\frac{\langle\Phi_{0}(x,z),\phi_{n}^{-}\rangle_{\mathcal{E}}}{D_{p}}\phi_{n}^{-}+\frac{\langle\Phi_{0}(x,z),\phi_{n}^{+}\rangle_{\mathcal{E}}}{D_{P}}\phi_{n}^{+}\right)\cos(\omega_{n}t)\,{\rm d}k\\ -\frac{1}{\rho}\sum_{n=0}^{N}\int_{0}^{\infty}\left(\frac{\langle P_{0}(x,z),\phi_{n}^{-}\rangle_{\mathcal{E}}}{D_{p}}\phi_{n}^{-}+\frac{\langle P_{0}(x,z),\phi_{n}^{+}\rangle_{\mathcal{E}}}{D_{p}}\phi_{n}^{+}\right)\frac{\sin(\omega_{n}t)}{\omega_{n}}\,{\rm d}k.$ (21)
4 Orthogonality of the eigenfunctions
Next, we determine the normalisation coefficients $D_{p}(k)$ while simultaneously confirming the orthogonality of the eigenfunctions. Firstly, we calculate the $L^{2}(-h,0)$ inner product of the depth-dependent functions, yielding
$\int_{-h}^{0}f_{n}(z;k)f_{p}(z;\tilde{k})\,{\rm d}z=\int_{-h}^{0}\frac{\displaystyle\cosh{\mu_{n}(z+h)}}{\cosh{\mu_{n}h}}\frac{\displaystyle\cosh{\tilde{\mu}_{p}(z+h)}}{\cosh{\tilde{\mu}_{p}h}}\,{\rm d}z$
$\hskip 101.17755pt=\frac{1}{\mu_{n}}\tanh{\mu_{n}h}-\frac{\tilde{\mu}_{p}}{\mu_{n}^{2}}\tanh{\tilde{\mu}_{p}h}+\frac{\tilde{\mu}_{p}^{2}}{\mu_{n}^{2}}\int_{-h}^{0}f_{n}(z;k)f_{p}(z;\tilde{k})\,{\rm d}z,$
[…]
$\int_{-h}^{0}f_{n}(z;k)f_{p}(z;\tilde{k})\,{\rm d}z=\frac{1}{g}\frac{\omega_{n}^{2}-\tilde{\omega}_{p}^{2}}{\mu_{n}^{2}-\tilde{\mu}_{p}^{2}},$ (22b)
[…]
$\quad\quad+\frac{c^{2}}{g}\int_{-\infty}^{\infty}\varphi_{n}^{\pm}(x,0;k)\overline{\varphi_{p}^{\pm}(x,0;\tilde{k})}\,{\rm d}x,$
[…]
where $\delta(\kappa)$ is the Dirac delta. Finally, in the case of $n=p$, we can recover the normalisation constant by taking limits,
$\lim_{k\rightarrow\tilde{k}}\langle\phi_{n}^{\pm}(\cdot,\cdot,k),\phi_{p}^{\pm}(\cdot,\cdot,\tilde{k})\rangle_{\mathcal{E}}=\frac{c^{2}}{g}\left(\frac{\tilde{k}}{\mu_{n}\partial_{k}\tilde{\mu}_{n}}\right)\delta(k-\tilde{k})\delta_{np},$
where
$\partial_{k}\mu_{n}=\frac{2\dfrac{c^{2}}{g}k}{\mu_{n}h{\rm sech}^{2}{\mu_{n}h}+\tanh{\mu_{n}h}+2\dfrac{c^{2}}{g}\mu_{n}}.$
For simplicity, we write
$\lim_{k\rightarrow\tilde{k}}\langle\phi_{n}^{\pm}(\cdot,\cdot,k),\phi_{p}^{\pm}(\cdot,\cdot,\tilde{k})\rangle_{\mathcal{E}}=D_{p}(k)\delta(k-\tilde{k})\delta_{np},$
where
$D_{p}(k)=\left(\frac{h}{2}{\rm sech}^{2}{\mu_{p}h}+\frac{\tanh{\mu_{p}h}}{2\mu_{p}}+\frac{c^{2}}{g}\right).$
Similarly,
$\langle\phi_{n}^{\pm},\phi_{p}^{\mp}\rangle_{\mathcal{E}}=\frac{c^{2}}{g}\left(\frac{k^{2}-\tilde{k}^{2}}{\mu_{n}^{2}-\tilde{\mu}_{p}^{2}}\right)\delta(k+\tilde{k})=0,$ (23c)
since $k$ and $\tilde{k}$ are both positive real numbers.
5 Time-domain simulation
We will now compute time-domain solutions of the internal pressure
$P(x,z,t)=-\rho\Phi_{t}(x,z,t),$ (24)
and free-surface
$\eta(x,t)=-\frac{1}{g}\Phi_{t}(x,0,t)=\frac{1}{\rho g}P(x,0,t).$ (25)
Noting that when computing the pressure over the full $z$-domain, we recover the surface as a byproduct.
5.1 Numerical evaluation
When evaluating (24), care must be taken due to the potentially highly oscillatory nature of the integrand (for large $x$ or $t$). We reformulate the integrals into the form
$I_{n}^{(1)}$ $=\sum_{\pm_{1}}\sum_{\pm_{2}}\pm_{2}\int_{0}^{\infty}g_{n\pm_{1}}^{(1)}e^{{\rm i}(\pm_{1}kx\pm_{2}\omega_{n}t)}\,{\rm d}k,$
$I_{n}^{(2)}$ $=\sum_{\pm_{1}}\sum_{\pm_{2}}\int_{0}^{\infty}g_{n\pm_{1}}^{(2)}e^{{\rm i}(\pm_{1}kx\pm_{2}\omega_{n}t)}\,{\rm d}k,$
where
$g_{n\pm}^{(1)}=\frac{\omega_{n}\langle\Phi_{0}(x,z),\phi_{n}^{\pm}\rangle_{\mathcal{E}}f_{n}(z)}{2D_{p}}\quad\text{and}\quad g_{n\pm}^{(2)}=\frac{\langle P_{0}(x,z),\phi_{n}^{\pm}\rangle_{\mathcal{E}}f_{n}(z)}{2D_{p}}.$
Here, the subscripts of $\pm$ indicate whether the signs are chosen independently (different subscripts) or must be consistent (same subscript). This formulation allows us to implement a Filon-type quadrature method (Dahlquist & Björck, 2008; Filon, 1930) detailed in Appendix A. Further efficiency can be gained by recognising that
$g_{+}e^{{\rm i}(kx+\omega_{n}t)}=\overline{g_{-}e^{-{\rm i}(kx+\omega_{n}t)}},$
halving the number of terms in the summations that must be computed.
5.2 Underwater pressure disturbance
As an example, we will consider an underwater pressure disturbance as the initial condition (generated by an underwater explosion, for example). The fluid is assumed initially at rest, $\Phi_{0}(x,z)=0$, with the initial pressure distribution
$P_{0}(x,z)=\hat{P}e^{\displaystyle-\pi^{2}\frac{(x-x_{c})^{2}+(z-z_{c})^{2}}{\sigma^{2}}},$ (26)
where $\hat{P}$ is the maximum pressure, $(x_{c},z_{c})$ is the centre of the pressure distribution and $\sigma$ is the size of the pressure. For the following results, we choose $\hat{P}=10^{6}$ Pa, $(x_{c},z_{c})=(0,-2000)$ m, and $\sigma=200$m.
Figure 2: Compression waves caused by an initial Gaussian pressure distribution centred at (0,2000)(0,-2000).
Figure 2 shows the solutions at $t=1$, 22, and $6.75$s. We can see that the initial pressure spreads radially away from the initial location at the speed of sound. The expanding wave reflects off both the bottom and the free surface, the phase of the pressure wave reversing as it reflects off the surface (Figure 2b). This process continues as time evolves, resulting in multiple pressure waves propagating away from the pressure origin. The effect this has on the free surface is the creation of wave crests above the initial pressure location that propagate away. The free-surface elevation between the crests will also be greater than at equilibrium. Full animation is given as Movie $1$.
6 Inclusion of static compression
In this section, we include (mild) density variance of the ocean water, the effect which is known in the literature as static compression, into the mathematical formulation, and develop the solution using a similar approach to that used in the previous section. The rigorous derivation of the governing equations can be found in the works of Longuet-Higgins (1950), Miyoshi (1954), and Sells (1965). Note that in the subsequent mathematical treatment, the density $\rho$ represents its value at the surface, which is treated as the baseline density. The hydrostatic density is $\rho e^{-\gamma z}$ (Abdolali & Kirby, 2017), where the effect of the variance in the density away from the surface is reflected through the term $\gamma$, which depends on the speed of sound in the water and is defined by $\gamma=g/c^{2}$. The effect of static compression is small for realistic oceans, as we shall see, and it significantly complicates the formulation. We believe that including it is worthwhile, first because it allows accurate calculation of its effect and second because it highlights how powerful the special inner product-based method is and how it can be extended to highly non-trivial cases. We note that the complexity arises in the theory, but that the computational difficulty is not greatly increased.
6.1 Mathematical Solution with Static Compression
The modified IBVP is given by
$\nabla^{2}\Phi$ $=\frac{1}{c^{2}}\Phi_{tt}+\gamma\Phi_{z},\quad$ $(x,z)\in\Omega,$ (27a)
$\Phi_{tt}+g\Phi_{z}$ $=0,\quad$ $z=0,$ (27b)
$\Phi_{z}$ $=0,$ $z=-h,$ (27c)
$\Phi_{x}$ $\to 0,$ $|x|\to\infty,$ (27d)
subject to the initial conditions
$\Phi(x,z,0)$ $=0,$ (28)
$\Phi_{t}(x,z,0)$ $=-\frac{P_{0}(x,z)}{\rho e^{-\gamma z}}.$ (29)
In order to identify an appropriate inner product, we apply Green’s second identity to the functions $\hat{\Phi}=e^{-\gamma z/2}\Phi$ and $\hat{\Psi}=e^{-\gamma z/2}\Psi$, where $\Phi$ and $\Psi$ satisfy (27), in the form
$\iint_{\Omega}\left(\nabla^{2}\hat{\Phi}\,\hat{\Psi}-\hat{\Phi}\,\nabla^{2}\hat{\Psi}\right){\rm d}V=\oint_{\partial\Omega}\left(\hat{\Phi}_{n}\,\hat{\Psi}-\hat{\Phi}\,\hat{\Psi}_{n}\right){\rm d}S,$
Taking $\Omega$ to be the region $(-L,L)\times(-h,0)$ gives, in the limit as $L\to\infty$
$\iint_{\Omega}e^{-\gamma z}\left[\left(\nabla^{2}\Phi-\gamma\Phi_{z}\right)\Psi-\Phi\left(\nabla^{2}\Psi-\gamma\Psi_{z}\right)\right]{\rm d}V=\int_{-\infty}^{\infty}\left.\left(\Phi_{z}\,\Psi-\Phi\,\Psi_{z}\right)\right|_{z=0}{\rm d}x,$
where we have used (27c) and (27d). Subsequently
$\iint_{\Omega}e^{-\gamma z}\left(\nabla^{2}\Phi-\gamma\Phi_{z}\right)\Psi{\rm d}V-\int_{-\infty}^{\infty}\left.\Phi_{z}\,\Psi\right|_{z=0}{\rm d}x=\iint_{\Omega}e^{-\gamma z}\Phi\left(\nabla^{2}\Psi-\gamma\Psi_{z}\right){\rm d}V$
$\hskip 216.81pt-\int_{-\infty}^{\infty}\left.\Phi\,\Psi_{z}\right|_{z=0}{\rm d}x.$ (30a)
Note that the combination of (27a) and (27b) yields $\Phi_{z}=-\frac{c^{2}}{g}(\nabla^{2}\Phi-\gamma\Phi_{z})$ on $z=0$, which further modifies the above expression into
$\iint_{\Omega}e^{-\gamma z}\mathcal{A}\Phi\Psi{\rm d}V+\frac{c^{2}}{g}\int_{-\infty}^{\infty}\left.\mathcal{A}\Phi\,\Psi\right|_{z=0}{\rm d}x$ $=\iint_{\Omega}e^{-\gamma z}\Phi\mathcal{A}\Psi{\rm d}V+\frac{c^{2}}{g}\int_{-\infty}^{\infty}\left.\Phi\,\mathcal{A}\Psi\right|_{z=0}{\rm d}x,$ (30b)
where $\mathcal{A}=\nabla^{2}-\gamma\partial_{z}$. Consequently, this operator is symmetric provided we choose the inner product to be
$\langle\Phi,\Psi\rangle_{\mathcal{E}}=\int_{\Omega}e^{-\gamma z}\Phi(x,z)\overline{\Psi(x,z)}\,{\rm d}x{\rm d}y+\frac{c^{2}}{g}\int_{-\infty}^{\infty}\Phi(x,0)\overline{\Psi(x,0)}\,{\rm d}x.$ (31)
In a way similar to what is mentioned in §3, we expand the time domain solution in the generalised eigenfunctions of the underlying differential operator. These eigenfunctions satisfy
$\nabla^{2}\phi$ $=-\frac{\omega^{2}}{c^{2}}\phi+\gamma\phi_{z},\quad$ $(x,z)\in\Omega,$ (32)
$-\omega^{2}\phi+g\phi_{z}$ $=0,\quad$ $z=0,$ (33)
$\phi_{z}$ $=0,$ $z=-h,$ (34)
and are of the form
$\phi_{n}^{\pm}(x,z;k)=e^{{\rm i}kx}f_{n}(z;k).$ (35)
In the case of static compression, the depth-dependent functions $f_{n}(z,k)$ are more complicated. They satisfy the eigenvalue problem
$f_{n}^{\prime\prime}(z;k)-\gamma f_{n}^{\prime}(z;k)$ $=\left(\mu_{n}^{2}-\frac{\gamma^{2}}{4}\right)f_{n}(z;k),$ (36a)
$f_{n}^{\prime}(-h;k)$ $=0,$ (36b)
$f^{\prime}(0;k)$ $=\frac{c^{2}}{g}\left(k^{2}+\frac{\gamma^{2}}{4}-\mu_{n}^{2}\right)f(0;k),$ (36c)
where the prime denotes differentiation with respect to $z$. For each $k$, the discrete spectrum of eigenvalues $\mu_{n}$ of (36) satisfy the following dispersion relation
$\frac{c^{2}}{g}\left(k^{2}+\frac{\gamma^{2}}{4}-\mu_{n}^{2}\right)\left(\displaystyle 1-\left(\frac{\gamma}{2\mu_{n}}\right)\tanh{\mu_{n}h}\right)=\mu_{n}\displaystyle\left[1-\left(\frac{\gamma}{2\mu_{n}}\right)^{2}\right]\tanh{\mu_{n}h}.$ (37)
The solutions to (37) are taken so that $\mu_{0}(k)\in\mathbb{R}$ and $\mu_{n}(k)$ are purely imaginary with positive and increasing imaginary part. The angular frequency associated with the mode $\phi_{n}^{\pm}(\cdot,\cdot;k)$ is
$\omega_{n}(k)=c\sqrt{k^{2}+\frac{\gamma^{2}}{4}-\mu_{n}^{2}.}$ (38)
Without loss of generality, we set $f(0)=1$, yielding
$f_{n}(z;k)=$ $\frac{\displaystyle e^{(\gamma z/2)}\left(\frac{\gamma}{2\mu_{n}}\sinh{\mu_{n}(z+h)}-\cosh{\mu_{n}(z+h)}\right)}{\frac{\gamma}{2\mu_{n}}\sinh{(\mu_{n}h)}-\cosh{(\mu_{n}h)}}.$ (39)
As in §3, the spectral theorem gives a time domain solution of the form
$\Phi(x,z,t)=\sum_{n=0}^{\infty}\int_{0}^{\infty}\left(A_{n}^{-}(k)\phi_{n}^{-}(x,z;k)+A_{n}^{+}(k)\phi_{n}^{+}(x,z;k)\right)\cos(\omega_{n}(k)t){\rm d}k\\ +\sum_{n=0}^{\infty}\int_{0}^{\infty}\left(B_{n}^{-}(k)\phi_{n}^{-}(x,z;k)+B_{n}^{+}(k)\phi_{n}^{+}(x,z;k)\right)\frac{\sin(\omega_{n}(k)t)}{\omega_{n}(k)}{\rm d}k,$ (40)
where the coefficient functions $A_{n}^{\pm}$ and $B_{n}^{\pm}$ are to be determined from the initial conditions. With orthonormality conditions again being of the form (14), albeit with different normalisation functions $D_{p}(k)$, the coefficients are recovered as
$A_{p}^{\pm}(k)=\frac{\langle\Phi_{0}(x,z),\phi_{p}^{\pm}(\cdot,\cdot,k)\rangle_{\mathcal{E}}}{D_{p}}\quad\mbox{and}\quad B_{p}^{\pm}(k)=-\frac{\langle P_{0}(x,z)/\rho e^{-\gamma z},\phi_{p}^{\pm}(\cdot,\cdot,k)\rangle_{\mathcal{E}}}{\displaystyle D_{p}},$ (41)
which is precisely in the form of (20). To finalise the mathematical solution in the case of static compression, it remains to compute the normalisation functions $D_{p}(k)$. This is done in Appendix B.
6.2 Time-domain simulations with static compression
Firstly, we consider the underwater explosion example detailed in §5.2, to compare with the results in the absence of static compression. We can see in Figure 3 (the full animation is in Movie $2$) that the results are visually indistinguishable from the prior results without static compression (Figure 2). To better observe the effects of static compression, we plot the percentage difference
$\frac{P_{\mathrm{static}}(x,z,t)-P(x,z,t)}{\max\limits_{x,z}(P_{\mathrm{static}}(x,z,t),P(x,z,t))},$
for $t=1$ (Figure 4), where $P_{\mathrm{static}}$ is the solution including static compression. We observe that the difference wave has a negative leading front in the upper half wave (i.e., upwards travelling waves) and a positive leading front in the lower half (i.e., downward travelling waves). We can see that the inclusion of static compression, on average, causes the magnitude of pressure above (below) the initial explosion location to decrease (increase) compared with the case with no static compression. It aligns with the fact that water is less (more) dense in the upper (lower) water region measured from the point of explosion when static compression is considered. There is a roughly $\pm 0.5\%$ pressure difference for the time considered ($0<t<20$) when including static compression. Finally, we note that the pressure wave speed has not changed with the addition of static compression. The Movie $3$ provides the full animation.
Figure 3: Compression waves caused by an initial Gaussian pressure distribution centred at (0,2000)(0,-2000), including the effects of static compression. In each panel, the line graph represents the free surface displacement, and the colour plot represents the hydrodynamic pressure distribution below the surface.
Figure 4: A plot of the pressure difference between Figure 2(a) and Figure 3(a) as a percentage of that maximum magnitude of pressure at that time ($t=1$), where a positive difference indicates that the pressure solution including static compression is greater than the pressure solution without static compression. The bottom panel provides a cross-section of the pressure difference at $x=0$.
We will now consider two different initial pressure distributions that have the same surface displacement but drastically different subsurface profiles. The first initial pressure distribution is the Gaussian distribution (26) centred at the free-surface $(x_{c},z_{c})=(0,0)$, and the second is a one-dimensional Gaussian with the same $x$ behaviour as (26),
$P_{1}(x,z)=\hat{P}\exp\left({\displaystyle-\pi^{2}\frac{(x-x_{c})^{2}}{\sigma^{2}}}\right).$ (42)
Comparing the results for the localised Gaussian (Figure 5) with those for the one-dimensional Gaussian (Figure 6), we can see that the initial pressures excite propagating gravity waves on the surface of the same shape, which move much slower than the pressure waves. These gravity waves are orders of magnitude greater than the surface displacement due to compression waves. This means that in the case that gravity waves and pressure waves are created by the same source, it would be hard to detect the effects of compression and further determine the initial pressure profile from surface observations. The pressure due to the gravity waves decays away from the surface, and therefore, underwater pressure measurements are better suited to observe the internal compression waves.
Focusing now on the internal pressure waves for our two initial pressures, as can be expected, the radically different initial pressure distributions lead to different internal pressure behaviour. The pressure wave generated by the localised Gaussian (Figure 5) propagates radially, reflecting off the bottom and surface, creating a similar surface knocking phenomenon to the submerged Gaussian. The one-dimensional Gaussian (Figure 6), on the other hand, predominately propagates in the $x$-direction with a rapidly decaying trail that reflects off the surface, which does not produce an observable knocking behaviour. The respective animations are provided in Movie $4$ and Movie $5$.
Figure 5: Compression waves caused by an initial Gaussian pressure distribution centred at $(0,0)$, including the effects of static compression.
Figure 6: Compression waves caused by an initial one-dimensional Gaussian pressure distribution centred at $x_{c}=0$, including the effects of static compression.
7 Conclusion
Under the assumptions of linearised water wave theory, this paper has considered the IBVP governing the evolution of a pressure disturbance in a two-dimensional ocean with a free surface, including the effect of compression. The problem was solved mathematically using self-adjoint operator theory by introducing an appropriate inner product. We presented two versions of the theory; in the first, we ignored the effect of static compression and included only the dynamic compression, while the second included both dynamic and static compression. Time-domain simulations showed the propagation of the pressure wave away from the initial source, reflecting off the free surface and the rigid ocean floor, and the onset of surface gravity waves propagating away from the source. In particular, a near reversal of phase is observed upon reflection from the surface, while the opposite pattern occurs in the reflection process from the rigid ocean floor. This indicates the fact that the free surface is approximately a homogeneous Dirichlet boundary. The corresponding time-domain simulations for the model incorporating static compression show that the effect of static compression is minor. While the present work deals with the canonical problem of uniform water depth, the theory considered here can be extended to tackle more realistic variable bathymetry using step-approximation, and it is equally applicable in modelling three-dimensional waves, albeit with increased computational requirements.
Supplementary material
Animations for Figures 26 are uploaded as supplementary material (https://doi.org/10.6084/m9.figshare.30186271.v1).
Funding
BW and MHM are supported by the Australian Research Council’s Discovery Projects funding scheme (DP240102104).
Declaration of interests
The authors report no conflict of interest.
Data Availability
Data sharing does not apply to this article, as no new data was created or analysed in this study. The computer code used to generate the solutions is available from the corresponding author upon reasonable request.
Appendix A Filon-type quadrature
Following the method of Filon (1930), we compute an integral of the form
$I=\int_{a}^{b}f(x)e^{{\rm i}\psi(x)}\,dx,$
where $f(x)$ and $\psi^{\prime}(x)$ are slowly varying functions, by splitting the domain of integration at the points $x_{0}=a,x_{1},\ldots,x_{N}=b$ and approximating the functions $f(x)$ and $\psi(x)$ by piecewise linear functions
$I$ $=\sum_{i=0}^{N-1}\int_{x_{i}}^{x_{i+1}}\left(f_{i}+\frac{f_{i+1}-f_{i}}{\Delta x_{i}}(x-x_{i})\right)e^{{\rm i}\left(\psi_{i}+\frac{\psi_{i+1}-\psi_{i}}{\Delta x_{i}}(x-x_{i})\right)}dx,$
$=\sum_{i=0}^{N-1}\Delta x_{i}\int_{0}^{1}\left(f_{i}+(f_{i+1}-f_{i})s\right)e^{{\rm i}\left(\psi_{i}+(\psi_{i+1}-\psi_{i})s\right)}ds,\quad\mbox{where }s=\frac{x-x_{i}}{\Delta x_{i}},$
$=\sum_{i=0}^{N-1}\Delta x_{i}\left[-\frac{{\rm i}}{\psi_{i+1}-\psi_{i}}\left(f_{i+1}e^{{\rm i}\psi_{i+1}}-f_{i}e^{{\rm i}\psi_{i}}\right)+\frac{f_{i+1}-f_{i}}{(\psi_{i+1}-\psi_{i})^{2}}\left(e^{{\rm i}\psi_{i+1}}-e^{{\rm i}\psi_{i}}\right)\right],$
$=\sum_{i=0}^{N-1}\Delta x_{i}\left[\left(\frac{{\rm i}e^{{\rm i}\psi_{i}}}{\psi_{i+1}-\psi_{i}}-\frac{e^{{\rm i}\psi_{i+1}}-e^{{\rm i}\psi_{i}}}{(\psi_{i+1}-\psi_{i})^{2}}\right)f_{i}+\left(-\frac{{\rm i}e^{{\rm i}\psi_{i+1}}}{\psi_{i+1}-\psi_{i}}+\frac{e^{{\rm i}\psi_{i+1}}-e^{{\rm i}\psi_{i}}}{(\psi_{i+1}-\psi_{i})^{2}}\right)f_{i+1}\right],$
$=\Delta x_{0}\left(\frac{{\rm i}e^{{\rm i}\psi_{0}}}{\psi_{1}-\psi_{0}}-\frac{e^{{\rm i}\psi_{1}}-e^{{\rm i}\psi_{0}}}{(\psi_{1}-\psi_{0})^{2}}\right)f_{0}+\sum_{i=1}^{N-1}\left[\Delta x_{i-1}\left(-\frac{{\rm i}e^{{\rm i}\psi_{i}}}{\psi_{i}-\psi_{i-1}}+\frac{e^{{\rm i}\psi_{i}}-e^{{\rm i}\psi_{i-1}}}{(\psi_{i}-\psi_{i-1})^{2}}\right)\right$.
$\quad\left.+\Delta x_{i}\left(\frac{{\rm i}e^{{\rm i}\psi_{i}}}{\psi_{i+1}-\psi_{i}}-\frac{e^{{\rm i}\psi_{i+1}}-e^{{\rm i}\psi_{i}}}{(\psi_{i+1}-\psi_{i})^{2}}\right)\right]f_{i}+\Delta x_{N-1}\left(-\frac{{\rm i}e^{{\rm i}\psi_{N}}}{\psi_{N}-\psi_{N-1}}+\frac{e^{{\rm i}\psi_{N}}-e^{{\rm i}\psi_{N-1}}}{(\psi_{N}-\psi_{N-1})^{2}}\right)f_{N}.$
The result is a weighting scheme for quadrature integration that can account for a rapidly changing phase function.
Appendix B Normalisation coefficients in the case of static compression
In this section, we compute the normalisation coefficients $D_{p}(k)$ needed in the solution in the case of static compression, as discussed in §6. We begin by introducing a new inner product over the vertical ordinate
$\langle u_{1},u_{2}\rangle_{h}=\int_{-h}^{0}e^{-\gamma z}u_{1}(z)\overline{u_{2}(z)}\,{\rm d}z+\frac{c^{2}}{g}u_{1}(0)\overline{u_{2}(0)},$ (43)
and seek to compute $\langle f_{n}(\cdot;k),f_{m}(\cdot;\tilde{k})\rangle_{h}$. Following some tedious calculations, we find
$\left(\mu_{n}^{2}-\frac{\gamma^{2}}{4}\right)\int_{-h}^{0}e^{-\gamma z}f_{n}(z;k)f_{m}(z;\tilde{k})\,{\rm d}z=\left[e^{-\gamma z}f_{n}^{\prime}(z;k)f_{m}(z;\tilde{k})\right]_{-h}^{0}-\left[e^{-\gamma z}f_{n}(z;k)f_{m}^{\prime}(z;\tilde{k})\right]_{-h}^{0}$
$\hskip 144.54pt+\int_{-h}^{0}e^{-\gamma z}f_{n}(z;k)\left(\tilde{\mu}_{m}^{2}-\frac{\gamma^{2}}{4}\right)f_{m}(z;\tilde{k})\,{\rm d}z,$ (44a)
which yields
$\int_{-h}^{0}e^{-\gamma z}f_{n}(z;k)f_{m}(z;\tilde{k})\,{\rm d}z=\frac{\omega_{n}^{2}(k)-\omega_{m}^{2}(\tilde{k})}{g\left(\mu_{n}^{2}-\tilde{\mu}_{m}^{2}\right)}.$ (44b)
Now, using (43),
$\langle f_{n}(z;k),f_{m}(z;\tilde{k})\rangle_{h}$ $=\frac{\omega_{n}^{2}(k)-\omega_{m}^{2}(\tilde{k})}{g\left(\mu_{n}^{2}-\tilde{\mu}_{m}^{2}\right)}f_{n}(0;k)f_{m}(0;\tilde{k})+\frac{c^{2}}{g},$
$=\frac{c^{2}}{g}\frac{k^{2}-\tilde{k}^{2}}{\mu_{n}^{2}-\tilde{\mu}_{m}^{2}}.$ (45a)
When $n=m$, the above inner product reduces to
$\lim_{k\rightarrow\tilde{k}}\langle f_{m}(z;k),f_{m}(z;\tilde{k})\rangle_{h}=\frac{c^{2}}{g}\frac{\tilde{k}}{\tilde{\mu}_{m}\partial_{k}\tilde{\mu}_{m}},$ (45b)
where
$\partial_{k}\mu_{n}=\frac{\displaystyle 2\frac{c^{2}}{g}\left(1-\frac{\gamma}{2}\frac{\tanh{\mu_{n}h}}{\mu_{n}}\right)k}{\mathcal{A}(\mu_{n},k)\tanh{\mu_{n}h}+\mathcal{B}(\mu_{n},k)\mathrm{sech}^{2}{\mu_{n}h}+2\dfrac{c^{2}}{g}\mu_{n}},$ (45c)
with $\mathcal{A}(\mu_{n},k)=\left(1-\frac{\gamma c^{2}}{2g}\right)+\left(\frac{\gamma}{2\mu_{n}}\right)^{2}-\frac{\gamma c^{2}}{2g}\left(\left(\frac{k}{\mu_{n}}\right)^{2}+\left(\frac{\gamma}{2\mu_{n}}\right)^{2}\right),$
$\mathcal{B}(\mu_{n},k)=\mu_{n}h\left[1-\left(\frac{\gamma}{2\mu_{n}}\right)^{2}+\frac{\gamma c^{2}}{2g}\left(\left(\frac{k}{\mu_{n}}\right)^{2}+\left(\frac{\gamma}{2\mu_{n}}\right)^{2}-1\right)\right].$
[…]
$\langle\phi_{n}^{\pm},\phi_{p}^{\pm}\rangle_{\mathcal{E}}$ $=\int_{-\infty}^{\infty}\int_{-h}^{0}\phi_{n}^{\pm}(x,z;k)\overline{\phi_{p}^{\pm}(x,z;\tilde{k})}\,{\rm d}z\,{\rm d}x+\frac{c^{2}}{g}\int_{-\infty}^{\infty}\phi_{n}^{\pm}(x,0;k)\overline{\phi_{p}^{\pm}(x,0;\tilde{k})}\,{\rm d}x,$
$=\int_{-\infty}^{\infty}\langle f_{n}(z;k),f_{p}(z;\tilde{k})\rangle_{h}e^{\mp{\rm i}(k-\tilde{k})x}\,{\rm d}x=\frac{c^{2}}{g}\frac{\tilde{k}}{\tilde{\mu}_{n}\partial_{k}\tilde{\mu}_{n}}\delta(k-\tilde{k})\delta_{np}.$ (46a)
Note that in the case of counter-propagating waves, we compute $\langle\phi_{n}^{\pm}(\cdot,\cdot,k),\phi_{p}^{\mp}(\cdot,\cdot,\tilde{k})\rangle_{\mathcal{E}}=0$, which is analogous to (23c).